OpenStudy (anonymous):

The line tangent to the curve y = 4 − x^2 at the point (2, 0) also passes through the point (0, a). Find a. Is a = 0?

OpenStudy (anonymous):

slope of the tangent = y' = -2x at (2,0) this slope = -4 equation of tangent is y-0 = -4(x-2) y = -4x + 8 when x = 0 y = 8 a = 8

OpenStudy (anonymous):

Thank you!